ten years from now, A will be twice as old as B. five years ago,A was three times as old as B. what are the present ages of A and B?

Answer by stanbon(75887)

ten years from now, A will be twice as old as B. five years ago,A was three times as old as B. what are the present ages of A and B?
------
Equations:
a+10 = 2(b+10)
a-5 = 3(b-5)
----------------
Rearrange:
a - 2b = 10
a - 3b = -10
----
Subtract and solve for "b":
b = 20 (B's age now)
----
Solve for "a" using a = 2b + 10
a = 2*20+10 = 50 (A's age now)
=====================
Cheers,
Stan H.
==============
-